# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ II.4 Multi-parameter estimation
$[\text{Cov}(\tilde{\boldsymbol{\theta}})]_{ij}=\langle\tilde{\theta}_{i}\tilde{\theta}_{j}\rangle-\langle\tilde{\theta}_{i}\rangle\langle\tilde{\theta}_{j}\rangle,$ (16)
@@
The matrix inequality in Eq. (15) means that $(\text{Cov}[\tilde{\boldsymbol{\theta}}]-I(\boldsymbol{\theta})^{-1})$ is a positive semidefinite matrix. While the above is the general case, dealing with this matrix form is challenging. Hence, it is convenient to transform the matrix inequality of Eq. (15) into a scalar bound by introducing a positive and real weight matrix $\mathcal{W}$ such that Albarelli et al. (2020):
$\text{Tr}[\mathcal{W}\text{Cov}[(\tilde{\boldsymbol{\theta}})]\geq\text{Tr}[\mathcal{W}I(\boldsymbol{\theta})^{-1}].$ (18)
@@
$[I_{Q}(\boldsymbol{\theta})]_{ij}=\frac{1}{2}\text{Tr}[\rho_{\boldsymbol{\theta}}\{L_{i},L_{j}\}].$ (20)
@@
$\text{Tr}[\mathcal{W}\text{Cov}[(\tilde{\boldsymbol{\theta}})]\geq\text{Tr}[\mathcal{W}I(\boldsymbol{\theta})^{-1}]\geq\text{Tr}[\mathcal{W}I_{Q}(\boldsymbol{\theta})^{-1}].$ (21)
@@
While the above tighter estimation bound includes the SLDs $L_{i}$ and $L_{j}$, other bounds can also be achieved by defining the Right Logarithmic Derivative Yuen and Lax (1973) and the Holevo approach Holevo (2011b, 1977). However, the precision bound using the SLDs already contains the same information found in both the Holevo and the RLDs bounds Albarelli et al. (2020). For this reason, it is common to use Eq. (19) with their corresponding matrix elements defined in Eq. (20) as the ultimate precision limit for the multi-parameter estimation case.
One of the main differences between the single- and multi-parameter scenarios is that, while the CFI can be achieved in both scenarios, in principle, the QFI is attainable only in the single-parameter case and not for the jointly multi-parameter estimation Albarelli et al. (2020). This can be understood because, generally, there is no single symmetric logarithmic derivative $L_{i}$ that defines an optimal measurement basis suitable for all the unknown parameters. To quantify this notion of incompatibility between parameters Ragy et al. (2016)—those that cannot be estimated optimally with a single set of measurements—the incompatibility matrix $\boldsymbol{D}$ has been introduced Albarelli et al. (2020); Carollo et al. (2019a), also known as mean Uhlmann curvature Carollo et al. (2019a), with elements:
$D_{ij}=-\frac{i}{2}\text{Tr}[\rho_{\boldsymbol{\theta}}[L_{i},L_{j}]].$ (23)
The case $D_{ij}=0$ for all $i,j$ is known as the compatibility condition Ragy et al. (2016). Recently, a genuine quantum incompatibility measure with a geometric character has also been proposed Belliardo and Giovannetti (2021), along with a comprehensive study of incompatibility under general $p$-local measurements Chen et al. (2022).
The expressions given by Eqs. (6)-(11) to obtain QFI and SLD for single parameter estimation, and generalizations thereof to the multiparameter case all involve diagonalizing the quantum state $\rho$ explicitly to find out its spectral decomposition. For quantum many-body density matrices whose dimensions increase exponentially with system-size, this generally presents a significant computational challenge. To overcome this, Ref. (Šafránek, 2018) noted that the Lyapunov equations can be expressed as a set of linear equations in the vectorization picture (Laub, 2005; Simoncini, 2016), and thereby provided the following formula for the QFI matrix
$[I_{Q}(\boldsymbol{\theta})]_{ij}=2\text{vec}[\partial_{\theta_{i}}\rho]^{\dagger}\left(\bar{\rho}\otimes I+I\otimes\rho\right)^{-1}\text{vec}[\partial_{\theta_{j}}\rho],$ (24)
@@
$\text{vec}[L_{i}]=2\left(\bar{\rho}\otimes I+I\otimes\rho\right)^{-1}\text{vec}[\partial_{\theta_{i}}\rho]$ (25)
@@ II.5 Geometry of parameter estimation
$I_{Q}(\boldsymbol{\theta})=4h_{\text{Bures}}=8\lim_{\delta\theta\rightarrow 0}\frac{1-F(\rho_{\theta},\rho_{\theta+\delta\theta})}{(\delta\theta)^{2}},$ (31)
@@ II.6 Alternatives to QFI for Metrology
Gauging the precision of parameter estimation through the quantum Cramér-Rao bound is undoubtedly the most popular approach in the literature, thanks to its geometrical properties and its utility as a signature of multipartite entanglement. However, the Cramér-Rao bound suffers from two practical difficulties. Firstly, although this bound is asymptotically tight for single parameter estimation, it may perform quite poorly in the non-asymptotic regime and especially if the likelihood function is highly non-Gaussian. This is a known issue even in classical parameter estimation theory, and several alternate bounds have been proposed in statistical literature (Van Trees and Bell, 2007). Perhaps the most well-known is the Ziv-Zakai bound (Ziv and Zakai, 1969), which is obtained by turning the continuous parameter estimation problem to a hypothesis testing problem for each coarse-grained interval in the domain of the parameter. By noting that hypothesis-testing at each interval amounts to a quantum state discrimination problem, Tsang in Ref. (Tsang, 2012) constructed a quantum version of the Ziv-Zakai bound. Although such bounds are not generally tight, they can be shown to perform significantly better for finite set of observations in highly non-Gaussian environments (Chang et al., 2022a). Extensions of the quantum Ziv-Zakai bound to multiparameter estimation (Berry et al., 2015) and noisy environments (Chang et al., 2022b) have also been investigated. Lu and Tsang also considered the quantum generalization of the classical Weiss-Weinstein bounds (Lu and Tsang, 2016). In yet another direction, Liu and Yuan (Liu and Yuan, 2016) obtained a tighter version of the Bayesian quantum Cramér-Rao bound valid for both biased and unbiased estimators. For a function $f(\theta)$ of the quantum parameter $\theta$ which is estimated via a (generally) biased estimator of bias $b(\theta)$ with a prior probability distribution $p(\theta)$ on a quantum state $\rho(\theta)$, Liu and Yuan’s bound reads
@@
where $I(\theta)$ is the QFI for $\rho(\theta)$ with respect to the parameter $\theta$. One can note that this bound reduces to the biased version of the Bayesian Quantum Cramér-Rao bound bound for a uniform prior. Secondly, and of significant relevance to many-body sensing, the QFI is not fully expressible in terms of experimentally easily accessible one and two-point correlation functions(Gabbrielli et al., 2018). This was addressed recently in Ref. (Lepori et al., 2023), where an alternative functional dependent only on one and two-point correlators was proposed. More economic sensing with fewer copies was studied in Ref. Zhang and Tong (2024).
@@ III Interferometric Quantum Sensing
The first notion of quantum sensing has been formulated in a seminal paper by Giovannetti, Lloyd and Maccone Giovannetti et al. (2004). In this case, the goal is to estimate the phase $\theta$ which is imprinted in quantum states through a unitary operation of the form $U(\theta)=e^{iH\theta}$, where $H$ is a Hermitian operator with real eigenvalues $E_{k}$ and their corresponding eigenvectors $|E_{k}\rangle$, such that $H|E_{k}\rangle=E_{k}|E_{k}\rangle$. This implies that the eigenvalues of the unitary operator $U(\theta)$ are given by $u_{k}=e^{iE_{k}\theta}$ while the eigenvectors do not depend on $\theta$ and remain the same as $H$, namely $U(\theta)|E_{k}\rangle=u_{k}|E_{k}\rangle$. The unknown parameter $\theta$ can be encoded in a quantum state through $|\Psi(\theta)\rangle=U(\theta)|\Psi_{0}\rangle$ for a given initial state $|\Psi_{0}\rangle$. In this case, one can easily show that the QFI of $|\Psi(\theta)\rangle$ is directly proportional to the variance of $H$
$I_{Q}(\theta)=4\left(\langle\Psi(\theta)|H^{2}|\Psi(\theta)\rangle-\langle\Psi(\theta)|H|\Psi(\theta)\rangle^{2}\right).$ (34)
One can maximize the above QFI with respect to the initial state. In fact, the maximum of $I_{Q}$ is achieved for the initial state $|\Psi_{0}\rangle=\left(|E_{\rm min}\rangle+|E_{\rm max}\rangle\right)/\sqrt{2}$, where $|E_{\rm min}\rangle$ ($|E_{\rm max}\rangle$) is the eigenvector of $H$ with the corresponding smallest (largest) eigenvalue $E_{\rm min}$ ($E_{\rm max}$). Thus, the corresponding QFI becomes
@@
To analyse the scaling of $I_{Q}$, in Eq. (35), with respect to the system size one has to consider the details of the interaction in the Hamiltonian $H$. For a given interaction of $H=\sum_{\{j_{1},\cdots,j_{k}\}}h_{j_{1}\cdots j_{k}}^{(k)}$ where $h_{j_{1}\cdots j_{k}}^{(k)}$ represents a $k$-body interaction between particles at sites $j_{1}$ to $j_{k}$, the QFI, at best, scales as $I_{Q}\sim N^{2k}$ Boixo et al. (2007), which is known as the Heisenberg limit. This shows enhanced precision in comparison with classical scenarios in which Fisher information, at best, scales linearly with $N$. The source of this enhancement is the special form of the superposition between the two energy eigenstates with extreme eigenvalues. As a special case, it is worth considering a probe of $N$ qubits which interact with a global magnetic field in the $z$ direction. In other words, $H=\sum_{j=1}^{N}\sigma_{j}^{z}$, with $\sigma_{j}^{z}$ being the $z$-Pauli operator which acts on qubit $j$. In this case, $|E_{\rm min}\rangle=|1,1,\cdots,1\rangle$ and $|E_{\rm max}\rangle=|0,0,\cdots,0\rangle$ and thus the optimal initial state will be the GHZ state $|\Psi_{0}\rangle=\left(|0,0,\cdots,0\rangle+|1,1,\cdots,1\rangle\right)/\sqrt{2}$ and, consequently, the QFI becomes $I_{Q}=N^{2}$.
Note that the quantum enhanced sensitivity with interferometric approach relies on creating GHZ-type entangled states or equivalently $N00N$ Mitchell et al. (2004) or squeezed states Ma et al. (2011); Maccone and Riccardi (2020). Preparation of these special forms of entangled states is not an easy task and is resource consuming. In fact, delocalization of quantum correlation is found to be essential for fast generation of metrologically useful entangled states Chu et al. (2024). In Ref. Chu et al. (2023), a strong metrological limit for the preparation time of such large-scale entanglement has been established through studying the Lieb-Robinson light cone.
The interferometric quantum phase sensing has been implemented experimentally Mitchell et al. (2004); Pezzé et al. (2007); Ono et al. (2013); Spagnolo et al. (2012); Okamoto et al. (2012, 2017). However, the scalability of this scenario might be challenging as: (i) initialization to a superposition between the two extreme eigenstates, namely $|E_{\rm min}\rangle+|E_{\rm max}\rangle$, is experimentally challenging; (ii) such superposition is highly susceptible to decoherence which transforms the superposition into a classical mixture with no quantum advantage for sensing; and (iii) it is only limited to a special class of unitary operations $U(\theta)=e^{iH\theta}$.
@@ IV Equilibrium Quantum Sensing: Second Order Quantum Phase Transition
It is interesting to consider a more general unitary operation $\tilde{U}(\theta)=e^{i(\theta H+\tilde{H})}$ for encoding the parameter $\theta$, as $|\tilde{\Psi}(\theta)\rangle=\tilde{U}(\theta)|\Psi_{0}\rangle$. For such an unitary operation, the sensing performance is again quantified by QFI of the parameter $\theta$ denoted as $\tilde{I}_{Q}(\theta)$. Note that as far as $[H,\tilde{H}]\neq 0$, both the eigenvectors and eigenvalues of $\tilde{U}(\theta)$ depend on $\theta$. In Ref. De Pasquale et al. (2013) it has been shown that for all choices of initials states $|\Psi_{0}\rangle$ and perturbing interactions $\tilde{H}$ we always have $I_{Q}(\theta)\geq\tilde{I}_{Q}(\theta)$. This means that in the interferometric phase sensing scenario, the presence of perturbing interaction $\tilde{H}$ is always destructive. For the case of sensing an external field, $\tilde{H}$ can represent interaction among the particles of the probe. In such cases, the interaction between the constituents of the probe can reduce the performance of the system Boixo et al. (2007); De Pasquale et al. (2013); Skotiniotis et al. (2015); Pang and Brun (2014).
@@
An alternative method to interferometric sensing is to harness the interaction between the particles in strongly correlated many-body systems. The interplay between different parts of the Hamiltonian can induce phase transition in the system, also known as quantum criticality. Many-body systems exhibiting quantum criticality have been identified as resources for quantum enhanced sensitivity. Recently, C. Hotter et al. proposed a unified protocol that combines critical and conventional approaches to metrology Hotter et al. (2024). A more fundamental limit to many-body quantum metrology was recently obtained in Ref. Abiuso et al. (2024). In this section, we review various forms of quantum criticality for sensing and show that a common feature among all these systems is the emergence of a sort of “gap-closing” in their energy spectrum. This is an important observation as it clearly shows that any phase transition which is accompanied by gap-closing is a potential resource for quantum sensing.
@@
As we have discussed in the previous section, despite the theoretical appeal, interferometry-based quantum sensing may face practical limitations. An alternative approach for achieving quantum-enhanced sensitivity is to exploit quantum criticality. The most conventional manifestation of quantum criticality is the second-order phase transition which describes a large set of phenomena. Second-order quantum phase transition occurs at zero temperature, where quantum fluctuations lead to spontaneous symmetry-breaking at the critical point. Interestingly, this opens up the possibility of quantum enhanced sensing near criticality (Zanardi and Paunković, 2006). Consider a many-body probe which is described by the Hamiltonian
$H=\theta H_{1}+H_{2},$ (36)
@@
with $H_{1}$ representing the action of the target parameter $\theta$ on the probe, and $H_{2}$ representing the internal physics of the probe itself. The interplay between $H_{1}$ and $H_{2}$ often leads to rich phase diagrams with the strength of the parameter $\theta$ determining the relevant phase. When unhindered by thermal fluctuations, a second order quantum phase transition may occur at a specific value of $\theta=\theta_{c}$, where the ground state of the system undergoes a global transformation. There are several common features in second-order quantum phase transitions: (i) spontaneous symmetry breaking; (ii) the emergence of a scale invariance behavior for all observables in the system; (iii) the appearance of a diverging length scale $\xi$ which behaves as $\xi{\sim}|\theta-\theta_{c}|^{-\nu}$, where $\nu$ is a critical exponent; and (iv) an anti-crossing between the ground and the first excited state whose energy gap vanishes at the transition point as the system sizes reaches thermodynamic limit. As we will discuss, a significant enhancement in sensitivity is often associated with the critical points in the phase diagram.
@@
At this point, a distinction should be made between two different approaches to many-body critical quantum sensing. In the first approach, the ground state of a many-body system, potentially at criticality, is used as the input state for an interferometry-based quantum sensing. The critical ground state is a delicate superposition of several states which can reveal significant variance with respect to certain observables and thus maximizes the QFI in an interferometry-based sensing, as given in Eq.(34). For instance, in Ref. (Ma et al., 2011) many-body spin squeezed states are used as inputs to interferometric phase estimation protocols. Atom-interferometry based quantum metrology has grown substantially in recent years, and their detailed description lies outside the scope of the present review. We direct the interested reader to Ref. (Pezze et al., 2018) for a comprehensive review. However, even for interferometric phase sensors with equilibrium probe states of genuinely many-body quantum systems, Frerot and Roscilde (Frérot and Roscilde, 2018) showed the possibility of quantum criticality enhanced sensing. They considered the problem of estimating a phase angle $\phi$ for the collective spin operator $J=\sum_{i}S_{i}$, such that the unitary $U=\exp(i\phi J)$ represents the action of the interferometer. By considering the ground state of a transverse field Ising chain as the input of an interferometer, they showed that at the critical point, the scaling of precision of estimation of $\phi$ for a $N$-atom transverse field Ising chain in $d$-dimensions is given by
${\delta\phi}^{2}\geq N^{-1-\frac{\eta-1}{d_{m}}},$ (37)

### v3

@@ II.4 Multi-parameter estimation
$[\text{Cov}[\tilde{\boldsymbol{\Theta}}]]_{ij}=\langle\tilde{\Theta}_{i}\tilde{\Theta}_{j}\rangle-\langle\tilde{\Theta}_{i}\rangle\langle\tilde{\Theta}_{j}\rangle,$ (16)
@@
The matrix inequality in Eq. (15) means that $(\text{Cov}[\tilde{\boldsymbol{\Theta}}]-I(\boldsymbol{\theta})^{-1})$ is a positive semidefinite matrix. While the above is the general case, dealing with this matrix form is challenging. Hence, it is convenient to transform the matrix inequality of Eq. (15) into a scalar bound by introducing a positive and real weight matrix $\mathcal{W}$ such that Albarelli et al. (2020a):
$\text{Tr}[\mathcal{W}\text{Cov}[\tilde{\boldsymbol{\Theta}}]]\geq\text{Tr}[\mathcal{W}I(\boldsymbol{\theta})^{-1}].$ (18)
@@
$[I_{Q}(\boldsymbol{\theta})]_{ij}=\frac{1}{2}\text{Tr}[\rho(\boldsymbol{\theta})\{\hat{L}_{i}(\boldsymbol{\theta}),\hat{L}_{j}(\boldsymbol{\theta})\}].$ (20)
@@
$\text{Tr}[\mathcal{W}\text{Cov}[\tilde{\boldsymbol{\Theta}}]]\geq\text{Tr}[\mathcal{W}I(\boldsymbol{\theta})^{-1}]\geq\text{Tr}[\mathcal{W}I_{Q}(\boldsymbol{\theta})^{-1}].$ (21)
@@
While the above tighter estimation bound includes the SLDs $\hat{L}_{i}(\boldsymbol{\theta})$ and $\hat{L}_{j}(\boldsymbol{\theta})$, other bounds can also be achieved by defining the Right Logarithmic Derivative Yuen and Lax (1973) and the Holevo approach Holevo (2011b, 1977). However, the precision bound using the SLDs already contains the same information found in both the Holevo and the RLDs bounds Albarelli et al. (2020a). For this reason, it is common to use Eq. (19) with their corresponding matrix elements defined in Eq. (20) as the ultimate precision limit for the multi-parameter estimation case.
One of the main differences between the single- and multi-parameter scenarios is that, while the CFI can be achieved in both scenarios, in principle, the QFI is attainable only in the single-parameter case and not for the jointly multi-parameter estimation Albarelli et al. (2020a). This can be understood because, generally, there is no single symmetric logarithmic derivative $\hat{L}_{i}(\boldsymbol{\theta})$ that defines an optimal measurement basis suitable for all the unknown parameters. To quantify this notion of incompatibility between parameters Ragy et al. (2016)—those that cannot be estimated optimally with a single set of measurements—the incompatibility matrix $\boldsymbol{D}$ has been introduced Albarelli et al. (2020a); Carollo et al. (2019a), also known as mean Uhlmann curvature Carollo et al. (2019a), with elements:
$D_{ij}=-\frac{i}{2}\text{Tr}[\rho(\boldsymbol{\theta})[\hat{L}_{i}(\boldsymbol{\theta}),\hat{L}_{j}(\boldsymbol{\theta})]],$ (23)
@@
The expressions given by Eqs. (6)-(11) to obtain QFI and SLD for single parameter estimation, and generalizations thereof to the multiparameter case all involve diagonalizing the quantum state $\rho$ explicitly to find out its spectral decomposition. For quantum many-body density matrices whose dimensions increase exponentially with system-size, this generally presents a significant computational challenge. To overcome this, Ref. (Šafránek, 2018) noted that the Lyapunov equations can be expressed as a set of linear equations in the vectorization picture (Laub, 2004; Simoncini, 2016), and thereby provided the following formula for the QFI matrix
$[I_{Q}(\boldsymbol{\theta})]_{ij}$ $=$ $2\text{vec}[\partial_{\theta_{i}}\rho(\boldsymbol{\theta})]^{\dagger}\Big{(}\overline{\rho(\boldsymbol{\theta})}\otimes I$ (24)
$+$ $I\otimes\rho(\boldsymbol{\theta})\Big{)}^{-1}\text{vec}[\partial_{\theta_{j}}\rho(\boldsymbol{\theta})],$
[…]
$\text{vec}[\hat{L}_{i}(\boldsymbol{\theta})]=2\left(\overline{\rho(\boldsymbol{\theta})}\otimes\hat{I}+\hat{I}\otimes\rho(\boldsymbol{\theta})\right)^{-1}\text{vec}[\partial_{\theta_{i}}\rho(\boldsymbol{\theta})]$ (25)
@@ II.5 Geometry of parameter estimation
$I_{Q}(\theta)=4h_{\text{Bures}}=8\lim_{\delta\theta\rightarrow 0}\frac{1-F(\rho(\theta),\rho(\theta+\delta\theta))}{(\delta\theta)^{2}},$ (31)
@@ II.6 Alternatives to QFI for metrology
Gauging the precision of parameter estimation through the quantum Cramér-Rao bound is undoubtedly the most popular approach in the literature, thanks to its geometrical properties and its utility as a signature of multipartite entanglement. However, the Cramér-Rao bound suffers from two practical difficulties. Firstly, although this bound is asymptotically tight for single parameter estimation, it may perform quite poorly in the non-asymptotic regime and especially if the likelihood function is highly non-Gaussian. This is a known issue even in classical parameter estimation theory, and several alternate bounds have been proposed in statistical literature (Van Trees and Bell, 2007). Perhaps the most well-known is the Ziv-Zakai bound (Ziv and Zakai, 1969), which is obtained by turning the continuous parameter estimation problem to a hypothesis testing problem for each coarse-grained interval in the domain of the parameter. By noting that hypothesis-testing at each interval amounts to a quantum state discrimination problem, Tsang in Ref. (Tsang, 2012) constructed a quantum version of the Ziv-Zakai bound. Although such bounds are not generally tight, they can be shown to perform significantly better for finite set of observations in highly non-Gaussian environments (Chang et al., 2022). Extensions of the quantum Ziv-Zakai bound to multiparameter estimation (Berry et al., 2015) and noisy environments (Chang et al., 2022) have also been investigated. Lu and Tsang also considered the quantum generalization of the classical Weiss-Weinstein bounds (Lu and Tsang, 2016). In yet another direction, Liu and Yuan (Liu and Yuan, 2016) obtained a tighter version of the Bayesian quantum Cramér-Rao bound valid for both biased and unbiased estimators. For a function $f(\theta)$ of the quantum parameter $\theta$ which is estimated via a (generally) biased estimator of bias $b(\theta)$ with a prior probability distribution $p(\theta)$ on a quantum state $\rho(\theta)$, Liu and Yuan’s bound reads
@@
where $I_{Q}(\theta)$ is the QFI for $\rho(\theta)$ with respect to the parameter $\theta$. One can note that this bound reduces to the biased version of the Bayesian Quantum Cramér-Rao bound for a uniform prior. Secondly, and of significant relevance to many-body sensing, the QFI is not fully expressible in terms of experimentally easily accessible one and two-point correlation functions(Gabbrielli et al., 2018). This was addressed recently in Ref. (Lepori et al., 2023), where an alternative functional dependent only on one and two-point correlators was proposed. More economic sensing with fewer copies was studied in Ref. Zhang and Tong (2024).
@@ III Interferometric Quantum Sensing
V. Giovannetti, S. Lloyd, and L. Maccone Giovannetti et al. (2004) addressed the problem of estimating the unknown phase $\theta$ imprinted in a quantum state through a unitary operation of the form $\hat{U}(\theta)=e^{i\hat{H}\theta}$, where $\hat{H}$ is a Hermitian operator with real eigenvalues $E_{k}$ and their corresponding eigenvectors $|E_{k}\rangle$, such that $\hat{H}|E_{k}\rangle=E_{k}|E_{k}\rangle$. This implies that the eigenvalues of the unitary operator $\hat{U}(\theta)$ are given by $u_{k}=e^{iE_{k}\theta}$ while the eigenvectors do not depend on $\theta$ and remain the same as $\hat{H}$, namely $\hat{U}(\theta)|E_{k}\rangle=u_{k}|E_{k}\rangle$. The unknown parameter $\theta$ can be encoded in a quantum state through $|\Psi(\theta)\rangle=\hat{U}(\theta)|\Psi_{0}\rangle$ for a given initial state $|\Psi_{0}\rangle$. In this case, one can easily show that the QFI of $|\Psi(\theta)\rangle$ is directly proportional to the variance of $\hat{H}$
$I_{Q}(\theta)=4\left(\langle\Psi(\theta)|\hat{H}^{2}|\Psi(\theta)\rangle-\langle\Psi(\theta)|\hat{H}|\Psi(\theta)\rangle^{2}\right).$ (34)
One can maximize the above QFI with respect to the initial state. In fact, the maximum of $I_{Q}$ is achieved for the initial state $|\Psi_{0}\rangle=\left(|E_{\rm min}\rangle+|E_{\rm max}\rangle\right)/\sqrt{2}$, where $|E_{\rm min}\rangle$ ($|E_{\rm max}\rangle$) is the eigenvector of $\hat{H}$ with the corresponding smallest (largest) eigenvalue $E_{\rm min}$ ($E_{\rm max}$). Thus, the corresponding QFI becomes
@@
To analyse the scaling of $I_{Q}$, in Eq. (35), with respect to the system size one has to consider the details of the interaction in the Hamiltonian $\hat{H}$. For a given interaction of $\hat{H}=\sum_{\{j_{1},\cdots,j_{k}\}}\hat{h}_{j_{1}\cdots j_{k}}^{(k)}$ where $\hat{h}_{j_{1}\cdots j_{k}}^{(k)}$ represents a $k$-body interaction between particles at sites $j_{1}$ to $j_{k}$, the QFI, at best, scales as $I_{Q}\sim N^{2k}$ Boixo et al. (2007), see e.g. Ref. Luis et al. (2016) and Ref. Beau and del Campo (2017) for its generalization to many-body open systems. In this review, we consider this quadratic scaling as Heisenberg limit, although, in the literature, there is not much agreement on the definition of the Heisenberg limit. For instance, sometimes $I_{Q}\sim N^{2k}$, for $k$-body interacting $\hat{H}$, may be defined as Heisenberg limit. Indeed, the most fundamental limit for interferometry-based sensing is given in Eqs. (34) and (35).
Interferometry-based quantum sensing, presented above, has clear advantages. First, the quantum-enhanced sensitivity, i.e. super-linear scaling of the QFI with respect to the system size, is achievable for all values of the unknown parameter $\theta$. Second, the measurement basis which achieves this quantum-enhanced precision is independent of the unknown parameter $\theta$ which implies that no prior information about the unknown parameter is needed to achieve quantum-enhanced precision. These two features clearly make interferometry-based quantum sensors very useful. Although, this approach has its own limitations too. The first limitation comes from the fact that the quantum-enhanced sensitivity with interferometric approach relies on creating GHZ-type entangled states or equivalently N00N Mitchell et al. (2004) or squeezed states Ma et al. (2011); Maccone and Riccardi (2020). Preparation of these special forms of entangled states is not an easy task and is resource consuming. In fact, delocalization of quantum correlation is found to be essential for fast generation of metrologically useful entangled states Chu et al. (2024). In Ref. Chu et al. (2023), a strong metrological limit for the preparation time of such large-scale entanglement has been established through studying the Lieb-Robinson light cone. The second limitation of interferometry-based sensing is that it is only applicable to the situations where the unknown parameter is encoded through a phase shit operator, namely $\hat{U}(\theta)$. This obviously does not represent the most general encoding method and thus not every quantum probe lies in this category. The third limitation is the spacial form of the initial state, namely the GHZ-type state $|\Psi_{0}\rangle=\left(|E_{\rm min}\rangle+|E_{\rm max}\rangle\right)/\sqrt{2}$ which are also known as Schrödinger cat states. These states are hard to implement in practice and, in general, prone to decoherence which transforms the superposition into a classical mixture with no quantum advantage for sensing. In fact, robust and scalable realization of GHZ states is an active area of research with several recent developments Pogorelov et al. (2021); Zhuang et al. (2021); Moses et al. (2023); Bluvstein et al. (2024); Bao et al. (2024); Cao et al. (2024). The interferometric quantum sensing has been implemented experimentally in various platforms Mitchell et al. (2004); Pezzé et al. (2007); Ono et al. (2013); Spagnolo et al. (2012); Okamoto et al. (2012, 2017).
@@ IV Equilibrium Quantum Sensing: Second Order Quantum Phase Transition
It is interesting to consider a more general unitary operation $\hat{U}^{\prime}(\theta)=e^{i(\theta\hat{H}+\hat{H}^{\prime}})$ for encoding the parameter $\theta$, as $|\Psi(\theta)\rangle=\hat{U}^{\prime}(\theta)|\Psi_{0}\rangle$. For such an unitary operation, the sensing performance is again quantified by QFI of the parameter $\theta$ denoted as $I^{\prime}_{Q}(\theta)$. Note that as far as $[\hat{H},\hat{H}^{\prime}]\neq 0$, both the eigenvectors and eigenvalues of $\hat{U}^{\prime}(\theta)$ depend on $\theta$. In Ref. De Pasquale et al. (2013) it has been shown that for all choices of initials states $|\Psi_{0}\rangle$ and perturbing interactions $\hat{H}^{\prime}$ we always have $I_{Q}(\theta)\geq I^{\prime}_{Q}(\theta)$. This means that in the interferometric phase sensing scenario, the presence of perturbing interaction $\hat{H}^{\prime}$ is always destructive. For the case of sensing an external field, $\hat{H}^{\prime}$ can represent interaction among the particles of the probe. In such cases, the interaction between the constituents of the probe can reduce the performance of the system Boixo et al. (2007); De Pasquale et al. (2013); Skotiniotis et al. (2015); Pang and Brun (2014).
@@
An alternative method to interferometric sensing is to harness the interaction between the particles in strongly correlated many-body systems. The interplay between different parts of the Hamiltonian can induce phase transition in the system, also known as quantum criticality. Many-body systems exhibiting quantum criticality have been identified as resources for quantum-enhanced sensitivity. Recently, C. Hotter et al. proposed a unified protocol that combines critical and conventional approaches to metrology Hotter et al. (2024). A more fundamental limit to many-body quantum metrology was recently obtained in Ref. Abiuso et al. (2025). In this section, we review various forms of quantum criticality for sensing and show that a common feature among all these systems is the emergence of a sort of “gap-closing” in their energy spectrum. This is an important observation as it clearly shows that any phase transition which is accompanied by gap-closing is a potential resource for quantum sensing.
@@
As we have discussed in the previous section, despite the theoretical appeal, interferometry-based quantum sensing may face practical limitations. An alternative approach for achieving quantum-enhanced sensitivity is to exploit quantum criticality. The most conventional manifestation of quantum criticality is the second-order phase transition which describes a large set of phenomena. Second-order quantum phase transition occurs at zero temperature, where quantum fluctuations lead to spontaneous symmetry-breaking at the critical point. Interestingly, this opens up the possibility of quantum-enhanced sensing near criticality (Zanardi and Paunković, 2006). Consider a many-body probe which is described by the Hamiltonian
$\hat{H}=\theta\hat{H}_{1}+\hat{H}_{2},$ (36)
@@
with $\hat{H}_{1}$ representing the action of the target parameter $\theta$ on the probe, and $\hat{H}_{2}$ representing the internal physics of the probe itself. The interplay between $\hat{H}_{1}$ and $\hat{H}_{2}$ often leads to rich phase diagrams with the strength of the parameter $\theta$ determining the relevant phase. When unhindered by thermal fluctuations, a second order quantum phase transition may occur at a specific value of $\theta=\theta_{c}$, where the ground state of the system undergoes a global transformation. There are several common features in second-order quantum phase transitions: (i) spontaneous symmetry breaking; (ii) the emergence of a scale invariance behavior for all observables in the system; (iii) the appearance of a diverging length scale $\xi$ which behaves as $\xi{\sim}|\theta-\theta_{c}|^{-\nu}$, where $\nu$ is a critical exponent; and (iv) an anti-crossing between the ground and the first excited state whose energy gap vanishes at the transition point as the system sizes reaches thermodynamic limit. As we will discuss, a significant enhancement in sensitivity is often associated with the critical points in the phase diagram.
@@
At this point, a distinction should be made between two different approaches to many-body critical quantum sensing. In the first approach, the ground state of a many-body system, potentially at criticality, is used as the input state for an interferometry-based quantum sensing. The critical ground state is a delicate superposition of several states which can reveal significant variance with respect to certain observables and thus maximizes the QFI in an interferometry-based sensing, as given in Eq.(34). For instance, in Ref. (Ma et al., 2011) many-body spin squeezed states are used as inputs to interferometric phase estimation protocols. Atom-interferometry based quantum metrology has grown substantially in recent years, and their detailed description lies outside the scope of the present review. We direct the interested reader to Ref. (Pezze et al., 2018) for a comprehensive review. However, even for interferometric phase sensors with equilibrium probe states of genuinely many-body quantum systems, Frerot and Roscilde (Frérot and Roscilde, 2018) showed the possibility of quantum criticality enhanced sensing. They considered the problem of estimating a phase angle $\theta$ for the collective spin operator $\hat{\bm{J}}=\sum_{i}\hat{\bm{S}}_{i}$, with $\hat{\bm{S}}_{i}=(\hat{S}^{x}_{i},\hat{S}^{y}_{i},\hat{S}^{z}_{i})$ being the vectorized spin operator at site $i$, such that the unitary $\hat{U}=\exp(i\theta\hat{J}^{x,y,z})$ represents the action of the interferometer along the relevant axis. By considering the ground state of a transverse field Ising chain as the input of an interferometer, they showed that at the critical point, the scaling of precision of estimation of $\theta$ for a $N$-atom transverse field Ising chain in $d$-dimensions is given by
${(\delta\theta)}^{2}\geq N^{-1-\frac{\eta-1}{d_{m}}},$ (37)